Two charged particles are 8cm apart.They are moved until the force between them triples.How far apart are they now?
1
Expert's answer
2019-10-11T10:38:09-0400

Force"(F)" is inversely proportional to "r^2"

So, "\frac{F_1}{F_2}=\frac{r_2^2}{r_1^2}"

As "F_2=3F_1"

Hence, "r_2=\frac{8}{\sqrt3}=4.6188cm"